# Kernel-based Koopman approximants for control: Flexible sampling, error analysis, and stability

### v1

@@ 4.1. Koopman approximants, flexible sampling, and error bounds
$\leq\sqrt{2N}(L_{g}+L_{G}R)\big{(}\max_{\ell\in[1:d]}\|U_{\ell}^{\dagger}\|\big{)}\varepsilon\cdot\big{\|}\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}(x)\big{\|}_{1},$
and it remains to bound $\|\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}(x)\|_{1}$. For this, fix $x\in\Omega$ and consider
$\|\mathbf{K}_{\mathcal{X}}^{-1}\textbf{k}_{\mathcal{X}}(x)\|_{1}=\sum_{i=1}^{d}\underbrace{|e_{i}^{\top}\mathbf{K}_{\mathcal{X}}^{-1}\textbf{k}_{\mathcal{X}}(x)|}_{=\pm e_{i}^{\top}\mathbf{K}_{\mathcal{X}}^{-1}\textbf{k}_{\mathcal{X}}(x)}=\bigg{(}\sum_{i=1}^{d}(-1)^{\ell_{i}}e_{i}\bigg{)}^{\top}\mathbf{K}_{\mathcal{X}}^{-1}\textbf{k}_{\mathcal{X}}(x),$
where $\ell_{i}\in\{0,1\}$ depends on $x$. Let $v_{x}=\sum_{i=1}^{d}(-1)^{\ell_{i}}e_{i}$ and set $h:=v_{x}^{\top}\mathbf{K}_{\mathcal{X}}^{-1}\textbf{k}_{\mathcal{X}}\in V_{\mathcal{X}}$. Then $h(x)=\|\mathbf{K}_{\mathcal{X}}^{-1}\textbf{k}_{\mathcal{X}}(x)\|_{1}$ for the fixed $x$, and thus, by Lemma 4.4, $\|\mathbf{K}_{\mathcal{X}}^{-1}\textbf{k}_{\mathcal{X}}(x)\|_{1}^{2}=h^{2}(x)\leq\mathsf{k}(x,x)\cdot v_{x}^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v_{x}$. Hence, noting that for the Wendland kernels it holds that $\mathsf{k}(x,x)=\Phi_{n,k}(0)$, we may estimate $\|\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}(x)\|_{1}^{2}\leq\Phi_{n,k}(0)\cdot\big{(}\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v\big{)}$, which shows $\big{|}H_{pq}(x)-\widehat{H}_{pq}(x)\big{|}\leq C\cdot D_{pq}(x)$ with
$D_{pq}(x)=h_{\mathcal{X}}^{k-1/2}$ $\operatorname{dist}(x,\mathcal{X})\|H_{pq}\|_{\mathbb{H}}$
$+\sqrt{2N}(L_{g}+L_{G}R)\Phi_{n,k}^{1/2}(0)\big{(}\max_{\ell\in[1:d]}\|U_{\ell}^{\dagger}\|\big{)}\big{(}\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v\big{)}^{1/2}\varepsilon.$
Finally, for $x\in\Omega$ and $u\in\mathbb{U}$ this gives
@@
$\leq$ $\max_{p}|H_{p,1}(x)-\widehat{H}_{p,1}(x)|+\max_{p}\big{|}\sum\nolimits_{q=2}^{m+1}\big{[}H_{pq}(x)-\widehat{H}_{pq}(x)]u_{q-1}\big{|}$
$\leq$ $C\max_{p}D_{p,1}(x)+C\max_{p,q}D_{p,q}(x)\sum_{q=1}^{m}|u_{q}|\leq C\max_{p,q}D_{p,q}(x)(1+\|u\|_{1})$
and the theorem is proved. ∎
Currently, the upper bound presented in Theorem 4.3 depends on $h_{\mathcal{X}}$, $N$ and $\varepsilon$. As indicated in Remark 4.2, for fixed $N$, more and more data points in the micro grid $\overline{\mathcal{X}}$ may be leveraged by using also a finer macro grid $\mathcal{X}$ (thus, decreasing $h_{\mathcal{X}}$) and decreasing the cluster radius $\varepsilon$, as more neighbors are contained in a smaller neighborhood.
We conclude this section by a remark concerning alternative control sampling schemes and the dependence of the error bound on the number cluster elements $N$.
@@
(a) If it is possible to sample the controlled flow map $f(x,u)$ at $(x_{\ell},u_{\ell,i})$, $\ell\in[1:d]$, $i\in[1:N]$, we can choose $x_{\ell,i}=x_{\ell}$ in Step 2 of Algorithm 1 and thus $\varepsilon=0$. Hence, we obtain the error bound
$\|f(x,u)-\widehat{f}(x,u)\|_{\infty}\leq C\cdot h_{\mathcal{X}}^{k-1/2}\operatorname{dist}(x,\mathcal{X})\max_{p,q}\|H_{pq}\|_{\mathbb{H}}(1+\|u\|_{1}).$
@@
(b) Let us briefly discuss the behavior of the term $\sqrt{N}\|U_{\ell}^{\dagger}\|$ in the estimate in Theorem 4.3 when the number of clusters $d$ remains constant and the cluster size $N$ grows. For this, we let $v_{i}:=\left[\begin{smallmatrix}1\\ u_{\ell,i}\end{smallmatrix}\right]\in\mathbb{R}^{m+1}$, $S:=\sum_{s=1}^{N}v_{i}v_{i}^{\top}$, and observe that $U_{\ell}U_{\ell}^{\top}=S$ and $U_{\ell}^{\dagger}=U_{\ell}^{\top}S^{-1}$. Therefore $\|U_{\ell}^{\dagger}\|^{2}=\|(U_{\ell}^{\dagger})^{\top}U_{\ell}^{\dagger}\|=\|S^{-1}U_{\ell}U_{\ell}^{\top}S^{-1}\|=\|S^{-1}\|=\frac{1}{\lambda_{\min}(S)}$, where $\lambda_{\min}(S)$ denotes the smallest eigenvalue of $S$. Hence, we have $\sqrt{N}\|U_{\ell}^{\dagger}\|=\sqrt{N/\lambda_{\min}(S)}$, and [45, Theorem 5.1.1] shows that if the $\bar{u}_{i}$ are drawn independently and uniformly from $[-R,R]^{m}$ and $R\geq\sqrt{3}$, then
@@
$\mathbb{P}\big{(}\sqrt{N}\|U_{\ell}^{\dagger}\|\geq c\big{)}\,\leq\,(1+m)\exp\big{(}-(1-\tfrac{1}{c})^{2}\cdot\tfrac{N}{1+R^{2}}\big{)},$
@@
hence it is exponentially unlikely that $\sqrt{N}\|U_{\ell}^{\dagger}\|$ is large if $N$ is large.
4.2. Feedback laws for stabilization
@@
Finally, we consider stabilizing feedback laws for the original control-affine system Equation 23 and its surrogate Equation 25, where $\widehat{f}$ is defined via Algorithm 1. As it turns out, under certain assumptions, a stabilizing feedback for Equation 25 is also stabilizing for Equation 23.
@@
Assume that the entries of $H=[g_{0}\;\;G]$ are contained in $\mathbb{H}$. Further, assume that there is a feedback law $\kappa:\Omega\to\mathbb{R}^{n}$, bounded on bounded sets [^8] One may also simply assume that the feedback law is admissible, i.e., $\kappa:\Omega\to\mathbb{U}$. such that $\widehat{F}_{\mathrm{cl}}(x):=\widehat{f}(x,\kappa(x))$ is asymptotically stable towards an equilibrium $x^{*}\in\Omega$ with $x^{*}\in\mathcal{X}$ with a Lyapunov function $V$ having a modulus of continuity $\omega_{V}$. Then $x^{*}$ is practically asymptotically stable w.r.t. $f(\cdot,\kappa(\cdot))=:F_{\rm cl}$, and the practical region decreases with decreasing fill distance $h_{\mathcal{X}}$ of the macro grid $\mathcal{X}$ and cluster size $\varepsilon$.
If further the samples are drawn in the sense of Remark 4.5(a) and if $\widehat{F}_{\mathrm{cl}}(x):=\widehat{f}(x,\kappa(x))$ is asymptotically stable towards $x^{*}\in\Omega$ with a Lyapunov function satisfying Assumption (a) or (b) of Theorem 3.9, then $\kappa$ also asymptotically stabilizes the original dynamics defined in Equation 23 towards $x^{*}$.
@@
We mimic the proofs of Theorem 3.5 and Theorem 3.9 for the closed loop system. In the first case, we observe that the error bound of Theorem 4.3 yields
@@
where we invoked the boundedness of $\kappa$ and with constants $c_{1}:=C\max_{p,q}\|H_{pq}\|_{\mathbb{H}}$, $c_{2}:=C\sqrt{2N}(L_{g}+L_{G}R)\Phi_{n,k}^{1/2}(0)\big{(}\max_{\ell\in[1:d]}\|U_{\ell}^{\dagger}\|\big{)}\big{(}\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v\big{)}^{1/2}$ and $c_{3}:=\max\{c_{1},c_{2}\}(1+\sup_{x\in\Omega}\|\kappa(x)\|_{1})$. This error bound is structurally the same as the bound of Corollary 2.8 with the cluster size $\varepsilon$ taking the role of $\sqrt{\lambda}$ and $\omega_{\Upsilon}=\operatorname{id}$. Thus, a similar argumentation as in the proof of Theorem 3.5 yields the claim.
To prove the second claim, we note that in view of the sampling of Remark 4.5(a), we may set $\varepsilon=0$ and obtain
$\leq h_{\mathcal{X}}^{k-1/2}\|x-x^{*}\|\max_{p,q}\|H_{pq}\|_{\mathbb{H}}(1+\sup_{x\in\Omega}\|\kappa(x)\|_{1}),$
@@
4.3. Numerics
Next, we illustrate the result of Theorem 4.3 by means of a numerical example. Therefore, consider the dynamics of the discretized Duffing oscillator given by
$x^{+}=F(x,u)=x+\Delta t\begin{bmatrix}x_{2}\\ x_{1}-3x_{1}^{3}u\end{bmatrix}=\underbrace{\begin{bmatrix}{x}_{1}+\Delta tx_{2}\\ {x}_{2}+\Delta tx_{1}\end{bmatrix}}_{:=g_{0}(x)}+\underbrace{\begin{bmatrix}0\\ -\Delta t3x_{1}^{3}\end{bmatrix}}_{:=G(x)}u$ (30)
with $\Delta t=0.05$ on $\Omega=[-2,2]^{2}$. As in Section 3.3, we use the coordinate functions $\psi_{i}(x)=x^{\top}e_{i}$, $i\in\{1,2\}$, as observables and for the Wendland radial basis functions, we choose the smoothness degree $k=1$. Further, whenever we refer to randomly drawn data samples, we always consider the uniform distribution on the respective set.
For the generation of the data points, we first use Chebyshev nodes to create the grid $\mathcal{X}\subset\Omega$ consisting of $d$ points that will act as the set of cluster points in the macro grid. With randomly chosen control values $u_{i}\in\mathbb{U}=[-2,2]$, the data triples $(x_{i},u_{i},x_{i}^{+})$ for $i\in[1:d]$ can be assembled.
To approximate the functions $g_{0}$ and $G$ at the grid points in $\mathcal{X}$ as presented in Step 1 of Algorithm 1, for a data point $x_{\ell}\in\mathcal{X}$ we choose the number of neighbors $N=25$ and then the neighbors $x_{\ell,i}$ for $i\in[1:N]$ are randomly drawn from a ball around $x_{\ell}$ with radius $\varepsilon=\frac{1}{d}$. For each $x_{\ell,i}$ a random $u_{\ell,i}\in\mathbb{U}$ is also drawn which yields the data points $(x_{\ell,i},u_{\ell,i},x^{+}_{\ell,i})$ for all $\ell\in[1:d]$ needed for the regression problem. Note that $\varepsilon$ is chosen to decease with an increasing number of cluster points. This is to make sure, that $\varepsilon$ can compensate the factor $\big{(}\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v\big{)}^{1/2}$ in the error estimate in Theorem 4.3 in view of the decreasing smallest eigenvalue of $\mathbf{K}_{\mathcal{X}}$ when decreasing the fill distance of the macro grid $\mathcal{X}$, that is, increasing $d$.
For Step 2, the approximated values of $g_{0}$ and $G$ on the macro grid $\mathcal{X}$ from Step 1 are used for the interpolation as stated in Algorithm 1 to obtain the control-affine surrogate system $\widehat{F}$ in (30).
For the first simulation, we construct a kEDMD-based surrogate $\widehat{F}$ using a macro grid of $d=6561$ cluster points. Let $x(\cdot;x^{0},u)$ and $\widehat{x}(\cdot;x^{0},u)$ denote the flow from the initial value $x^{0}\in\Omega$ with control $u$ for the original model $F$ and surrogate $\widehat{F}$, respectively. In Figure 4, we inspect the difference of these trajectories for the initial condition $x^{0}=\binom{0.1}{0.1}$ and control sequences $\{u_{i}\}$ of length 50, where ten control values are randomly chosen from $\mathbb{U}$ and each of these control values is applied for five time steps (cf. Figure 4 on the right). In the phase space plot in the left of Figure 4, the two trajectories are barely distinguishable. When comparing the absolute error in the middle of Figure 4 on a log scale, we observe that the surrogate maintains an absolute error less than $10^{-4}$ for approximately 40 time steps (corresponding to the time $t=2$ in the continuous model).

### v3

@@ 4.1. Koopman approximants, flexible sampling, and error bounds
$\leq\sqrt{2N}(L_{g}+L_{G}R)\big(\max\nolimits_{\ell\in[1:d]}\|U_{\ell}^{\dagger}\|\big)\varepsilon\cdot\big\|\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}(x)\big\|_{1}.$
To bound the last term, fix $x\in\Omega$ and
$\|\mathbf{K}_{\mathcal{X}}^{-1}\textbf{k}_{\mathcal{X}}(x)\|_{1}=\sum\nolimits_{i=1}^{d}{|e_{i}^{\top}\mathbf{K}_{\mathcal{X}}^{-1}\textbf{k}_{\mathcal{X}}(x)|}=\sum\nolimits_{i=1}^{d}(-1)^{\ell_{i}}e_{i}^{\top}\mathbf{K}_{\mathcal{X}}^{-1}\textbf{k}_{\mathcal{X}}(x),$
where $\ell_{i}\in\{0,1\}$ depends on $x$. Let $v_{x}=\sum_{i=1}^{d}(-1)^{\ell_{i}}e_{i}$ and set $h:=v_{x}^{\top}\mathbf{K}_{\mathcal{X}}^{-1}\textbf{k}_{\mathcal{X}}\in V_{\mathcal{X}}$. Then $h(x)=\|\mathbf{K}_{\mathcal{X}}^{-1}\textbf{k}_{\mathcal{X}}(x)\|_{1}$ for the fixed $x$, and thus, by lemma 4.3, $\|\mathbf{K}_{\mathcal{X}}^{-1}\textbf{k}_{\mathcal{X}}(x)\|_{1}^{2}=h^{2}(x)\leq\mathsf{k}(x,x)\cdot v_{x}^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v_{x}$. Hence, noting that for the Wendland kernels it holds that $\mathsf{k}(x,x)=\Phi_{n,k}(0)$, we may estimate $\|\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}(x)\|_{1}^{2}\leq\Phi_{n,k}(0)\cdot\big(\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v\big)$, which shows $\big|H_{pq}(x)-\widehat{H}_{pq}(x)\big|\leq C\cdot D_{pq}(x)$, where $D_{pq}(x)=c\Phi_{n,k}^{1/2}(0)(\max\nolimits_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v)^{1/2}\cdot\max\nolimits_{\ell\in[1:d]}\|U_{\ell}^{\dagger}\|\varepsilon$ ++ $h_{\mathcal{X}}^{k-1/2}\operatorname{dist}(x,\mathcal{X})\|H_{pq}\|_{\mathbb{H}}$ with $c=\sqrt{2N}(L_{g}+L_{G}R)$. Finally, for $x\in\Omega$ and $u\in\mathbb{U}$ the claim follows from
@@
$\leq$ $\max\nolimits_{p}|H_{p,1}(x)-\widehat{H}_{p,1}(x)|+\max\nolimits_{p}\big|\sum\nolimits_{q=2}^{m+1}\big[H_{pq}(x)-\widehat{H}_{pq}(x)]u_{q-1}\big|$
$\leq$ $C\max\nolimits_{p}D_{p,1}(x)+C\max\nolimits_{p,q}D_{p,q}(x)\sum\nolimits_{q=1}^{m}|u_{q}|\leq C\max\nolimits_{p,q}D_{p,q}(x)(1+\|u\|_{1}).$
Note that the full rank condition on the controls may be viewed as a condition ensuring persistently exciting inputs as commonly used in data-driven control, see e.g. [10].
Currently, the upper bound presented in theorem 4.2 depends on $h_{\mathcal{X}}$, $N$, and $\varepsilon$. As indicated in the discussion following algorithm 1, for fixed $N$, more data points in the micro grid $\overline{\mathcal{X}}$ may be leveraged by using also a finer macro grid $\mathcal{X}$ (thus, decreasing $h_{\mathcal{X}}$) and decreasing the cluster radius $\varepsilon$, as more neighbors are contained in a smaller neighborhood.
If one choose choose sufficiently many control samples at each data point in algorithm 1, we may choose $\varepsilon=0$ and obtain the following corollary as a direct consequence of theorem 4.2.
Corollary 4.4.
Let the assumptions of theorem 4.2 hold, assume that we have samples of $f(x,u)$ at $(x_{\ell},u_{\ell,i})$, $\ell\in[1:d]$, $i\in[1:N]$ and set $x_{\ell,i}=x_{\ell}$ in Step 2 of algorithm 1. Then there are $C,h_{0}>0$ such that for $h_{\mathcal{X}}<h_{0}$ and any $(x,u)\in\Omega\times\mathbb{U}$,
$\|f(x,u)-\widehat{f}(x,u)\|_{\infty}\leq C\cdot h_{\mathcal{X}}^{k-1/2}\operatorname{dist}(x,\mathcal{X})\max\nolimits_{p,q}\|H_{pq}\|_{\mathbb{H}}(1+\|u\|_{1}).$
We conclude this section by a remark concerning the dependence of the error bound on the number cluster elements $N$ and a possible regularization.
@@
(a) Let us briefly discuss the behavior of the term $\sqrt{N}\|U_{\ell}^{\dagger}\|$ in the estimate in theorem 4.2 when the number of clusters $d$ remains constant and the cluster size $N$ grows. For this, we let $v_{i}:=\left[\begin{smallmatrix}1\\ u_{\ell,i}\end{smallmatrix}\right]\in\mathbb{R}^{m+1}$, $Y:=\sum_{s=1}^{N}v_{i}v_{i}^{\top}$, and observe that $U_{\ell}U_{\ell}^{\top}=Y$ and $U_{\ell}^{\dagger}=U_{\ell}^{\top}Y^{-1}$. Therefore $\|U_{\ell}^{\dagger}\|^{2}=\|(U_{\ell}^{\dagger})^{\top}U_{\ell}^{\dagger}\|=\|Y^{-1}U_{\ell}U_{\ell}^{\top}Y^{-1}\|=\|Y^{-1}\|=\frac{1}{\lambda_{\min}(Y)}$, where $\lambda_{\min}(Y)$ denotes the smallest eigenvalue of $Y$. Hence, we have $\sqrt{N}\|U_{\ell}^{\dagger}\|=\sqrt{N/\lambda_{\min}(Y)}$. In what follows, we will show that if the $u_{i}$ are drawn independently and uniformly from $R,R]^{m}$ and $c>\max\{1,\frac{\sqrt{3}}{R}\}$, then
@@
$\mathbb{P}\big(\sqrt{N}\|U_{\ell}^{\dagger}\|\geq c\big)\,\leq\,(m+1)\exp\left[-\tfrac{N\min\{1,\frac{R^{2}}{3}\}}{1+mR^{2}}\cdot\tfrac{\tau-\log\tau-1}{\tau}\right],$ (29)
@@
where $\tau:=c^{2}\min\{1,\frac{R^{2}}{3}\}>1$. Hence, it is exponentially unlikely that $\sqrt{N}\|U_{\ell}^{\dagger}\|$ is large if $N$ is large. To prove (29), let us set $X_{i}=v_{i}v_{i}^{\top}$. Then $\|X_{i}\|=\|v_{i}\|^{2}=1+\|u_{i}\|^{2}\leq 1+mR^{2}=:L$ and $\mathbb{E}[X_{i}]=\operatorname{diag}(1,\mathbb{E}[u_{i}u_{i}^{\top}])=\operatorname{diag}(1,\frac{R^{2}}{3}I_{m})$, so that $\mu_{\min}:=\lambda_{\min}(\mathbb{E}[Y])=N\lambda_{\min}(\mathbb{E}[X_{i}])=N\min\{1,\frac{R^{2}}{3}\}$. We now have
$\mathbb{P}\big(\sqrt{N}\|U_{\ell}^{\dagger}\|\geq c\big)=\mathbb{P}\big(\lambda_{\min}(Y)\leq N/c^{2}\big)=\mathbb{P}\Big(\lambda_{\min}(Y)\leq\delta\mu_{\min}\Big),$
where $\delta=1/\tau<1$. Finally, [54, Theorem 5.1.1] provides the following upper bound of the above probability, which proves (29):
$(m+1)\left[e^{-(1-\delta)-\delta\log\delta}\right]^{\nicefrac{{\mu_{\min}}}{{L}}}=(m+1)\exp\left(-\tfrac{\mu_{\min}}{L}\cdot\tfrac{\frac{1}{\delta}-\log\frac{1}{\delta}-1}{\frac{1}{\delta}}\right).$
(b) To alleviate a bad conditioning of the system eq. 27 for small fill distances, one may, in addition to the clustering procedure, consider the regularized kernel matrix analogously to eq. 7 at the cost of an additional error term proportional to the regularization parameter.
4.2. Numerical example: Open-loop error bound
We illustrate the result of theorem 4.2 by means of a numerical example. Therefore, consider the dynamics obtained by an explicit Euler discretization of a controlled Duffing oscillator given by
$x^{+}=f(x,u)=\begin{bmatrix}{x}_{1}+\Delta tx_{2}\\ {x}_{2}+\Delta tx_{1}\end{bmatrix}+\begin{bmatrix}0\\ -\Delta t3x_{1}^{3}\end{bmatrix}:=g_{0}(x)+G(x)u$ (30)
with $\Delta t=0.05$ on $\Omega=[-2,2]^{2}$. As in Section 3.3, we use the coordinate functions $\psi_{i}(x)=x^{\top}e_{i}$, $i\in\{1,2\}$, as observables and for the Wendland radial basis functions, we choose the smoothness degree $k=1$. Further, whenever we refer to randomly drawn data samples, we consider the uniform distribution on the respective set. For the generation of the data points, we use the Padua points introduced in Section 3.3 to create the grid $\mathcal{X}\subset\Omega$ consisting of $d$ points that will act as the set of cluster points in the macro grid. With randomly chosen control values $u_{i}\in\mathbb{U}=[-2,2]$, the data triples $(x_{i},u_{i},x_{i}^{+})$ for $i\in[1:d]$ satisfying the assumption of Algorithm 1 can be assembled.
[…]
For the first simulation, we construct a kEDMD-based surrogate $\widehat{F}$ using a macro grid of d=6555d={6555} cluster points. Let $x(\cdot;x^{0},u)$ and $\widehat{x}(\cdot;x^{0},u)$ denote the flow from the initial value $x^{0}\in\Omega$ with control $u$ for the original model $F$ and surrogate $\widehat{F}$, respectively. In fig. 3, we inspect the difference of these trajectories for the initial condition $x^{0}=\binom{0.1}{0.1}$ and control sequences $\{u_{i}\}$ of length 50, where ten control values are randomly chosen from $\mathbb{U}$ and each of these control values is applied for five time steps (cf. fig. 3 on the right). In the phase space plot in the left of fig. 3, the two trajectories are barely distinguishable. When comparing the absolute error in the middle of fig. 3 on a log scale, the surrogate maintains an absolute error less than $10^{-3}$ for nearly 50 time steps (corresponding to time $t=2$ in the continuous model).
Figure 3. Trajectories $x(\cdot;x^{0},u)$ of system eq. 30 and its kEDMD surrogate $\widehat{x}(\cdot;x^{0},u)$ generated with algorithm 1 (left) and their deviation $\|x(k;x^{0},u)-\widehat{x}(k;x^{0},u)\|_{2}$, $k\in[1:50]$, in norm w.r.t. a threshold $10^{-{3}}$ (middle) for initial value $x^{0}=(0.1,0.1)^{\top}$ for the control sequence $u$ (right).
Next, we randomly chose 20 initial values $\hat{x}_{i}$ in $0.5,0.5]^{2}\subset\Omega$, as well as 20 random input sequences $\{u_{i}\}\subset[-0.1,0.1]\subset\Omega$ of length 30. These smaller boxes ensure that the trajectories remain in the domain $\Omega$.
Finally, in fig. 4, we inspect the error of a one-step prediction in the Euclidean norm, similar to fig. 1 in Section 3.3. The learning of the kEDMD-based system is again performed for the three numbers of cluster points  $d\in\{435,1653,6555\}$ for the Padua-based macro grid and $N=25$ data points in a neighborhood of $\varepsilon=\nicefrac{{1}}{{d}}$. The errors $\max\nolimits_{j\in[1:20]}\|F(x,u_{j})-\widehat{F}(x,u_{j})\|$ on a respective validation grid in $1,1]^{2}$ are computed for the control values in $\{-2+0.2(j-1)|\ j\in[1:20]\}$. In the intensity plots of fig. 4, it can be observed that the error evenly decreases the greater $d$ is chosen. This validates our results from theorem 4.2: There are two sources of error that can be influenced, i.e., the fill distance of the macro grid $h_{\mathcal{X}}$ depending on number of cluster points $d$ and the maximal cluster size $\varepsilon$. If $\varepsilon$ is small enough to compensate the factor $\big(\max\nolimits_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v\big)^{\nicefrac{{1}}{{2}}}$ that increases with $d$, the decrease of the fill distance by refining $\mathcal{X}$ yields a decay of the approximation error.
Figure 4. Approximation error $\max_{j\in[1:20]}\|f(x,u_{j})-\widehat{f}(x,u_{j})\|$ for the predicted system eq. 30 with $d\in\{435,1653,6555\}$ cluster points using a Padua-based grid (from left to right).
4.3. Feedback stabilization
@@
Finally, we consider stabilizing feedback laws for the original control-affine system eq. 22 and its surrogate eq. 24, where $\widehat{f}$ is defined via algorithm 1. As it turns out, under certain assumptions, a stabilizing feedback for eq. 24 is also stabilizing for eq. 22.
@@
Let $H=[g_{0}\;\;G]$ be componentwise in $C_{b}^{\lceil\sigma_{n,k}\rceil}(\Omega)$ with $\sigma_{n,k}=(n+1)/2+k$ and $k\geq 1$. Further, let $\kappa:\Omega\to\mathbb{R}^{n}$ be a locally Lipschitz-continuous feedback law asymptotically stabilizing the equilibrium $x^{*}\in\Omega\cap\mathcal{X}$ of the closed-loop dynamics $\widehat{F}_{\mathrm{cl}}(x):=\widehat{f}(x,\kappa(x))$ such that the modulus of continuity $\omega_{V}$ of the respective Lyapunov function $V$ and is uniform in $h_{\mathcal{X}}\in(0,h_{0}]$ for some $h_{0}>0$. Then, $x^{*}$ is practically asymptotically stable w.r.t. $f(\cdot,\kappa(\cdot))=:F_{\rm cl}$ with practical region decreasing in the fill distance $h_{\mathcal{X}}$ of the macro grid $\mathcal{X}$ and the cluster size $\varepsilon$.
If further the samples are drawn with cluster size $\varepsilon=0$ and if $\widehat{F}_{\mathrm{cl}}(x):=\widehat{f}(x,\kappa(x))$ is asymptotically stable towards $x^{*}\in\Omega$ with a Lyapunov function satisfying Assumption (a) or (b) of theorem 3.9, then $\kappa$ also asymptotically stabilizes the original dynamics defined in eq. 22 towards $x^{*}$.
@@
We mimic the proofs of theorem 3.5 and theorem 3.9 for the closed-loop system. In the first case, we observe that the error bound of theorem 4.2 yields
@@
where we invoked the boundedness of $\kappa$ and with constants $c_{1}:=C\max_{p,q}\|H_{pq}\|_{\mathbb{H}}$, $c_{2}:=C\sqrt{2N}(L_{g}+L_{G}R)\Phi_{n,k}^{1/2}(0)\big(\max_{\ell\in[1:d]}\|U_{\ell}^{\dagger}\|\big)\big(\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v\big)^{1/2}$ and $c_{3}:=\max\{c_{1},c_{2}\}(1+\sup_{x\in\Omega}\|\kappa(x)\|_{1})$. This error bound is structurally the same as the bound of corollary 2.8 with the cluster size $\varepsilon$ taking the role of $\sqrt{\lambda}$ and $\omega_{\Upsilon}=\operatorname{id}$. Thus, we obtain a similar estimate to (17) with $\bar{C}h_{\mathcal{X}}^{1/2}$ replaced by $c_{3}(h_{\mathcal{X}}^{k+1/2}+\varepsilon)$. To mimic remainder of the proof of theorem 3.5, we remain to show the Lipschitz continuity of $\widehat{F}_{\mathrm{cl}}(x)=\widehat{g}_{0}+\widehat{G}(x)\kappa(x)$, uniformly in the fill distance $h_{\mathcal{X}}$. With the same argumentation as in lemma 3.6, we may first show that $\widehat{g}_{0}$ and $\widehat{G}$ are Lipschitz continuous. As the feedback is assumed to be locally Lipschitz, and as all functions are bounded due to continuity, $\widehat{F}_{\mathrm{cl}}$ is also locally Lipschitz continuous as sum and product of bounded and locally Lipschitz continuous functions. Thus, a similar argumentation as in the second part of the proof of theorem 3.5 yields the claim.
To prove the second claim, we note that in view of the sampling of corollary 4.4, we may set $\varepsilon=0$ and obtain
$\|F_{\mathrm{cl}}(x)-\widehat{F}_{\mathrm{cl}}(x)\|$ $\leq h_{\mathcal{X}}^{k-1/2}\|x-x^{*}\|\max\nolimits_{p,q}\|H_{pq}\|_{\mathbb{H}}(1+\sup\nolimits_{x\in\Omega}\|\kappa(x)\|_{1}),$
@@
Exemplarily, we consider the discrete-time system from [21, Example 3.1], i.e.,
$\begin{bmatrix}x_{1}^{+}\\ x_{2}^{+}\end{bmatrix}=\tilde{f}(x,u)=\begin{bmatrix}\exp(0.3x_{2})\sqrt{1+x_{1}+x_{2}}-0.4x_{2}-1\\ 0.5\ln(1+x_{1}+x_{2})+0.4x_{2}\end{bmatrix}+\begin{bmatrix}0.5\\ 0\end{bmatrix}u$ (31)
with map $\tilde{f}:\tilde{\Omega}\rightarrow\mathbb{R}^{2}$ and domain $\tilde{\Omega}=[-0.4,1]^{2}$. For vanishing input $u=0$, the origin $x^{*}=0$ clearly is an equilibrium of system (31). However, as shown in [21], this equilibrium is not asymptotically stable, as the linearization at the origin has an eigenvalue of modulus larger than one. To stabilize the origin, [21] introduces the stabilizing state feedback $\kappa(x)=-\ln(1+x_{1}+x_{2})$. Using this example, we illustrate that the converse statement of Proposition 4.6, i.e., if the original system (31) is asymptotically stable towards the origin for the feedback law $\kappa$, the origin is practically asymptotically stable w.r.t. the kEDMD-based surrogate system.
To this end, we approximate the system (31) using Algorithm 1. The set of virtual observation points $\mathcal{X}$ is chosen as a grid of $d\in\{435,1653,6555,11476\}$ Padua points, as introduced in Section 3.3 and the cluster radius $\varepsilon$ is again set to $\tfrac{1}{d}$. Then, for each point in $\mathcal{X}$, we randomly draw 25 data points $(x_{\ell,i},u_{\ell,i},x_{\ell,i}^{+})$ for $i\in[1:d]$ and $\ell\in[1:25]$ such that the assumptions of Algorithm 1 are satisfied. As before, the coordinate functions are chosen as observables and the Wendland kernel with smoothness degree $k=1$ is used. The results are depicted in Figure 5. For reasons of clarity, the trajectories are only shown for $d=435$ (left), as the phase portrait does not show noticeable differences for higher numbers of virtual observation points. The errors (mid, right), however, are presented for all $d\in\{435,1653,6555,11476\}$.